OpenStudy (anonymous):

3.Solve x2 + 18 = –9x. (1 point) 4.Solve 4x2 = 12x please help by solving completing the square please show work

OpenStudy (anonymous):

3. x=-2 4.x=3/4

OpenStudy (anonymous):

i cant see 3 clearly is it 2x or 2

OpenStudy (anonymous):

xsquared

OpenStudy (anonymous):

x^2

OpenStudy (anonymous):

3. x^2+18=-9x x^2+9x+18=0 (x+6)(x+3) Either x=-3 or x=-6 4. 4x^2=12x 4x^2-12x=0 4x(x-3)=0 Either x=0 or x=3

OpenStudy (anonymous):

x^2+9x+18=0 x^2+6x+3x+18=0 2x(x+3)+3(x+3) (2x+3)(x+3) x=-3,-3/2

OpenStudy (anonymous):

8/12=x x=3/4

OpenStudy (anonymous):

4.Solve 4x2 = 12x + 40. (1 point) typo this is the equation for #4

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

ok 8-40=2x -32=2x -32/2=x x=-16

OpenStudy (anonymous):

if is correct tell me

OpenStudy (anonymous):

4x^2=12x+40 4x^2-12x-40=0 let f(x)=4x^2-12x-40 (function) Replace the value of x in such a way to obtain zero f(-2)=0 (x+2) is a factor of f(x) 4x^2-12x-40=(x+2)(4x-20)=0 Either x=-2 or (4x-20=0>> x=5)

OpenStudy (anonymous):

thank yuh thunderemperor

OpenStudy (anonymous):

You're welcome. ;)